# Building an Open AIBOM Standard in the Wild

### v1

@@ 7.2. Road Ahead
Evolving the AIBOM to Capture FM-Powered Software. The rapid mainstream adoption of foundation models (FMs) and FM-powered software (FMware) since mid-2023 has fundamentally changed what an AIBOM specification must capture. FMware represents a clear departure from earlier paradigms. Whereas traditional AI-powered systems, using simple machine learning models such as Support Vector Machines or Neural Networks, are largely static artifacts with well-defined data sources and predictable lifecycle boundaries; FMware is continuous, adaptive, and open-ended (Hassan et al., 2024). It introduces new asset types such as prompts and downstream fine-tuned variants. Agents, which orchestrate workflows and mediate decision-making across multi-model pipelines, become first-class entities whose roles and interactions require explicit documentation (Hassan et al., 2024; Rajbahadur et al., 2024). FMware architectures also add operational layers such as grounding (linking decisions to verifiable sources) and guarding (enforcing safety, policy, and compliance), each producing additional artifacts and dependencies that must be tracked (Rajbahadur et al., 2024). These shifts, coupled with increasing regulatory expectations and enterprise risk-management requirements, make the evolution of the AIBOM to capture FMwareBOM both necessary and urgent. This next iteration expands beyond static system representation to capture agent behaviors, orchestration context, grounding sources, guardrail policies, data lineage, and lifecycle transformations. While our current AIBOM can adequately capture traditional AI-powered software, as evidenced by our validation in Section 5, we need to evolve it urgently for FMware.
@@
To address FMware’s expanded requirements, we are preparing an update to AI and Dataset profiles for an SPDX 3.1 release candidate (SPDX Working Group, 2025a) that introduces fields for agent identity and capabilities (SPDX Working Group, 2025c), richer provenance links connecting datasets to generation prompts and retrieval contexts (SPDX Working Group, 2025d), and mechanisms to capture adaptation events, fine-tuning lineage, and downstream modifications (SPDX Working Group, 2025b). These changes aim to ensure that FMwareBOM captures the full lifecycle of modern AI systems while remaining automatable, auditable, and compatible with evolving compliance practices.
@@
Making the Future AIBOM Tool-Native and Automatable. As highlighted in our Section 7.1, tooling is essential for adoption, and as discussed above, FMware significantly increases the scale and complexity of what must be tracked. To address this, we are designing the evolved version of AIBOM to be tooling-native from the outset. Our goal is to ensure that key fields can be extracted directly from development pipelines or inferred through lightweight analysis, while still meeting regulatory and risk-management requirements. Our hope is that this approach will enable greater automation throughout the FMware lifecycle—from generating BOMs and validating schema conformance during CI stages to assessing compliance and policy obligations automatically. Together, these design choices aim to make AIBOM not only richer in representation but also a practical, automatable, and auditable part of the FMware development process.
@@
Enabling Interoperability Across the AI Supply-Chain Ecosystem. A key focus of our roadmap is ensuring that the next generation of SPDX AIBOM does not evolve in isolation but becomes a connective layer across a fragmented ecosystem of complementary standards and initiatives. Currently, efforts such as SLSA (for provenance and build integrity) (OpenSSF / SLSA Project, 2025), OpenChain (for organisational compliance) (OpenChain Project / The Linux Foundation, 2025), Croissant (et al., 2024) and emerging security frameworks like Coalition for Secure AI (Coalition for Secure AI, 2024) all address critical but distinct layers of the AI supply chain, yet lack a unified integration point. To address this, we are actively working to align and interoperate with these complementary initiatives. For example, we are exploring how FMwareBOM can serve as the canonical machine-readable artifact for AI compliance within the OpenChain AI WG, collaborating with OpenSSF and CISA to integrate provenance, verification, and security metadata, and engaging directly with industry and open-source leaders through collaborative initiatives. Towards this goal, we organized a panel at Open Source Summit North America 2025 on harmonizing SLSA provenance and SPDX SBOM (Maruseac et al., 2025), soliciting community input on how the two standards can complement and strengthen each other. Our goal is to reduce duplication, enable seamless metadata flow across layers, and build a cohesive ecosystem where complementary standards reinforce one another. The SPDX 3.0 specification is being prepared for formal submission to ISO as part of the effort to update the ISO/IEC 5962 standard. This planned ISO standardization, which will involve reviews by member bodies and Technical Committees, is anticipated to significantly boost adoption and ecosystem growth.
@@ 8. Related Work
Emerging approaches for trustworthy AI. Multiple initiatives aim to improve AI system trustworthiness by enhancing auditability, traceability, compliance, and provenance. DataBOM treats datasets as first-class artifacts to enable provenance tracking (Liu et al., 2024), but provides limited model-level detail. CycloneDX ML-BOM extends SBOM practices to ML models and datasets (bac, 2024), yet overlooks lifecycle complexities. Structured documentation approaches like IBM AI Factsheets (bac, 2025), Model Cards (bac, 2022), and Datasheets (Gebru et al., 2021; Mitchell et al., 2019) improve reporting but lack standard compliance and holistic AI system level coverage. AI Cards (Golpayegani et al., 2024) capture richer risk information but require tooling support, while TAIBOM (NquiringMinds LTD, 2025) automates BOM creation with novel, ecosystem-incompatible schemas. Broader governance efforts, including the NIST AI RMF (AI, 2023), Microsoft’s Responsible AI Standard (Microsoft, 2022), and Croissant (et al., 2024), promote accountability but remain disconnected from software supply chain realities, with Croissant focusing primarily on dataset metadata.
Despite their contributions, these solutions remain fragmented, either treating artifacts like datasets or models in isolation or proposing high-level frameworks detached from implementation. This piecemeal landscape leaves practitioners without a single, machine-readable artifact offering an end-to-end view of an AI system’s composition and provenance. Our goal with AIBOM is to address this gap: by extending the dependency-aware structure of SBOMs, it unifies data, models, and lifecycle processes into a cohesive, auditable standard.
@@
SBOM research in SE. SBOMs are widely used for vulnerability management, transparency, risk assessment, and supply chain integrity (Qvarfordt, 2024; Martini, 2023). Empirical work shows that they reduce remediation times by enabling rapid dependency analysis (Benedetti et al., 2024) and improve auditability through provenance and licensing metadata (Martini, 2023). However, their impact depends heavily on tooling completeness and metadata quality (Stalnaker et al., 2024; Xia et al., 2023a). Researchers and practitioners alike have proposed extensions, such as blockchain-enabled SBOMs for tamper resistance (Xia et al., 2024) and DataBOMs for capturing provenance in data pipelines (Barclay et al., 2019; Liu et al., 2024). However, our study is the first document the process of creating AIBOM specification in a global, multi-stake holder setting.
Standardization Experience Reports. Existing experience reports on standardization in SE primarily offer historical narratives rather than structured, reproducible approaches to extension. Classic retrospectives on POSIX trace its evolution from user-group initiatives to IEEE and ISO ratification, providing valuable lessons but limited methodological insight into how such extensions were conceived and executed (Isaak, 2005; Jespersen, 1995). Complementary procedural documents, such as the Austin Group’s SD/6, detail governance aspects like defect handling and new work item procedures but remain detached from practitioner perspectives (Josey, 2012). Similar gaps exist in other domains: for example, Accellera’s SystemVerilog report documents a successful upstream-to-IEEE transition yet does not codify the development methodology (Sutherland, 2002). Our work complements and extends this literature by providing a detailed, transparent account of the end-to-end experience of developing an AI-focused extension, capturing decision-making, iteration, validation, and stakeholder engagement and maps standards development as an AR cycle that can serve as a replicable blueprint for future standardization efforts.
Multi-stakeholder studies in SE. Software engineering is fundamentally a socio-technical endeavor, shaped by the interactions of diverse actors including developers, managers, customers, and regulators. Multi-stakeholder studies investigate this complexity, analyzing the conflicting goals and socio-technical dependencies that influence project outcomes. Prior work highlights the benefits of incorporating multiple perspectives for improving requirements elicitation (Davis, 1993) and risk management (Alqahtani et al., 2017), particularly in large-scale projects (Hoda et al., 2017).
However, this literature also documents significant challenges, such as balancing heterogeneous interests and coordinating communication across organizational boundaries (Ralph, 2016). These challenges are magnified in the context of open, community-driven efforts, such as standards development, which represent a distinct and highly complex form of multi-stakeholder collaboration. Unlike projects within or between firms, creating a standard involves a decentralized group of volunteers with no single organizational authority, making consensus-building and governance particularly challenging. Therefore, the goal of this paper was to document the high-complexity domain of multi-stakeholder collaboration using open standard development as a unique case study.
Action Research in SE. Given the complex realities of software development, AR has emerged as a key methodology for structuring industry–academia collaboration, enabling researchers to intervene in real-world settings and iteratively refine tools and processes (Avison et al., 1999; Petersen et al., 2014). Its maturity in SE is reflected in recent efforts to formalize methodological rigor (Staron, 2025) and hybridize approaches, such as Design Science Action Research (DSAR), which combine artifact construction with organizational change (de Castro et al., 2025).
However, prior work shows that AR is typically confined to single organizations or bilateral partnerships, focusing on internal process improvement (Dittrich et al., 2024). Its application to open, community-governed initiatives producing public standards remains largely unexplored. To our knowledge, this work is the first to apply AR in such a setting, adapting it from a tool for organizational change to a framework for coordinating a global, multi-stakeholder community around a shared standardization goal.
@@ 9. Conclusion
Our experience demonstrates that Action Research can serve as a powerful framework for guiding standardization efforts in fast-moving, multi-stakeholder domains. By structuring the development of AIBOM specification around iterative diagnose-design-evaluate-reflect cycles, we showed how open collaboration, continuous feedback, and evidence-based iteration can transform fragmented community efforts into a coherent, widely applicable standard. The lessons distilled from this process provide a blueprint for future initiatives like our evolution of AIBOM specification in SPDX 3.1. Looking ahead, we envision AR-driven approaches playing a central role in shaping agile, trustworthy standards that evolve alongside the technologies they aim to govern.
Acknowledgements. The participation of Arthit Suriyawongkul in the WG is made possible through the financial support of Taighde Éireann – Research Ireland under Grant number 18/CRT/6224 (Research Ireland Centre for Research Training in Digitally-Enhanced Reality (d-real)). We express our sincere gratitude to Helen Oakley and Rhea Michael Anthony for detailing the validation they conducted with our AIBOM framework in their industrial setting, as reported in Section 5.4. Furthermore, we would like to thanks Prof. Daniel M. German and Prof. Zhen Ming (Jack) Jiang for their valuable feedback and mentoring. Finally, we would also like to acknowledge the valuable contribution of all the members who participated in our SPDX AI and Dataset Working Group meetings.

### v2

@@ 7.2. Road Ahead
Evolving AIBOM for FM-powered software. The 2023 surge in foundation models (FMs) and FM-powered software (FMware) necessitated an evolution of the AIBOM specification. Unlike static ML artefacts, FMware is continuous, adaptive, and open-ended (Hassan et al., 2024), introducing new assets like prompts, fine-tuned variants, and agents—the latter acting as first-class entities in multi-model pipelines (Hassan et al., 2024; Rajbahadur et al., 2024). Operational layers such as grounding (decision source verification) and guarding (safety/policy enforcement) generate further dependencies requiring tracking (Rajbahadur et al., 2024). Driven by regulatory and risk-management needs, the FMwareBOM iteration expands AIBOM to capture agent behaviours, orchestration context, and lifecycle transformations. While the current specification handles traditional AI (see Section 5), adapting it for FMware is now urgent.
@@
To meet FMware’s requirements, we are updating the AI and Dataset profiles for the SPDX 3.1 release candidate (SPDX contributors, 2025a). This update introduces fields for agent identity and capabilities (SPDX contributors, 2025c), enhanced provenance linking datasets to prompts and retrieval contexts (SPDX contributors, 2025d), and mechanisms to track adaptation events and fine-tuning lineage (SPDX contributors, 2025b). These enhancements ensure FMwareBOM captures the full AI lifecycle while remaining automatable, auditable, and compliant with evolving standards.
@@
Making the future AIBOM tool-native and automatable. As Section 7.1 notes, effective tooling is vital for adoption, particularly as FMware increases tracking scale and complexity. To address this, we are designing the evolved version of AIBOM to be tooling-native from the outset. Our goal is to ensure that key fields can be extracted directly from development pipelines or inferred through lightweight analysis, while still meeting regulatory and risk-management requirements. Our hope is that this approach will enable greater automation throughout the FMware lifecycle—from generating BOMs and validating schema conformance during CI stages to assessing compliance and policy obligations automatically. These design choices aim to make AIBOM not only richer in representation but also a practical, automatable, and auditable part of the FMware development process.
@@
Enabling interoperability across the AI supply-chain ecosystem. Our roadmap positions the next-generation SPDX AIBOM as a connective layer across a fragmented ecosystem. Rather than evolving in isolation, SPDX aims to provide a unified integration point for complementary standards—including SLSA (build provenance) (Project, 2025b), OpenChain (organisational compliance) (Project, 2025a), Croissant (data provenance) (et al., 2024), and Coalition for Secure AI (Coalition for Secure AI, 2024)—ensuring interoperability across the AI supply chain. To address this, we are actively working to align and interoperate with these initiatives. For example, we are exploring how FMwareBOM can serve as the canonical machine-readable artefact for AI compliance within the OpenChain AI WG and collaborating with OpenSSF and CISA to integrate provenance, verification, and security metadata. To this end, our Open Source Summit North America 2025 panel on harmonising SLSA and SPDX (Maruseac et al., 2025) gathered community input on reducing metadata duplication and ensuring standard reinforcement. Concurrently, the formal SPDX 3.0 specification is being prepared to update the ISO/IEC 5962 standard. This standardisation process, involving Technical Committee reviews, is expected to accelerate global adoption and ecosystem maturity.
@@ 8. Related Work
Emerging approaches for trustworthy AI. Various initiatives address AI trustworthiness through auditability and provenance, yet gaps remain. DataBOM (Liu et al., 2024) focuses on dataset provenance but lacks model-level detail, while CycloneDX ML-BOM (CycloneDX, 2024) extends SBOMs to ML without capturing full lifecycle complexities. Reporting frameworks like factsheets (et al., 2019), model cards (Face, 2022), and datasheets (Gebru et al., 2021; Mitchell et al., 2019) enhance documentation but lack standardisation and system-level coverage. AI Cards (Golpayegani et al., 2024) offer rich risk data but lack tooling. Broader governance frameworks (NIST AI RMF (of Standards and Technology, 2023), Microsoft RAI (Microsoft, 2022), and Croissant (et al., 2024)) promote accountability but remain decoupled from software supply chain realities. Recently, TAIBOM (TAIBOM, 2025) aims to provide formal component and dependency descriptions. Despite their contributions, these solutions remain fragmented, treating datasets or models in isolation or offering high-level frameworks detached from implementation. This piecemeal landscape lacks a machine-readable artifact providing an end-to-end view of AI system composition. AIBOM addresses this gap by extending the dependency-aware structure of SBOMs to unify data, models, and lifecycle processes into a cohesive, auditable standard.
@@
SBOM research in SE. SBOMs are widely used for vulnerability management, transparency, risk assessment, and supply chain integrity (Qvarfordt, 2024; Martini, 2023). Empirical work shows that they reduce remediation times by enabling rapid dependency analysis (Benedetti et al., 2024) and improve auditability through provenance and licensing metadata (Martini, 2023). However, their impact depends heavily on tooling completeness and metadata quality (Stalnaker et al., 2024; Xia et al., 2023b). Researchers and practitioners alike have proposed extensions, such as blockchain-enabled SBOMs for tamper resistance (Xia et al., 2024) and DataBOMs for capturing provenance in data pipelines (Barclay et al., 2019; Liu et al., 2024). However, our study is the first document the process of creating AIBOM specification in a global, multi-stake holder setting.
Standardisation experience reports. SE standardisation reports often provide historical narratives rather than reproducible extension methodologies. POSIX retrospectives trace evolution from user groups to ISO ratification but lack detail on how extensions were conceived (Isaak, 2005; Jespersen, 1995). Similarly, the Austin Group’s SD/6 details governance but lacks practitioner perspectives (Josey, 2012), while SystemVerilog reports document IEEE transitions without codifying development methods (Sutherland, 2002). We extend this literature by detailing the end-to-end development of an AI extension—capturing decision-making and validation—and mapping standards development as a replicable Action Research (AR) cycle.
Multi-stakeholder studies in SE. Software engineering’s socio-technical nature requires balancing the conflicting goals of developers, managers, customers, and regulators. Research shows that analysing these multi-stakeholder dependencies improves requirements elicitation (Davis, 1993) and risk management (Alqahtani et al., 2017), particularly within large-scale agile environments (Hoda et al., 2017). Challenges like interest balancing and cross-organisational communication (Ralph, 2016) are magnified in decentralised standards development. Lacking a single organisational authority, these volunteer-driven efforts face unique consensus-building and governance hurdles. We therefore use open standard development as a case study to analyse high-complexity multi-stakeholder collaboration.
Action Research in SE. AR has emerged as a key methodology for industry–academia collaboration, enabling iterative refinement of tools in real-world settings (Avison et al., 1999; Petersen et al., 2014). While recent guidelines (Staron, 2025) and hybrid models like Design Science Action Research (de Castro et al., 2025) have matured the field, AR remains largely limited to internal process improvement within single organisations (Dittrich et al., 2024). We extend AR to open, community-governed initiatives, using it as a framework to coordinate global stakeholders around a shared standardisation goal.
@@ 9. Conclusion
Our experience establishes Action Research as a robust framework for standardisation in dynamic, multi-stakeholder domains. By structuring the development of AIBOM specification around iterative diagnose-design-evaluate-reflect cycles, we showed how open collaboration, continuous feedback, and evidence-based iteration can transform fragmented community efforts into a coherent, widely applicable standard. The lessons distilled from this process provide a blueprint for future initiatives like our evolution of AIBOM specification in SPDX 3.1. Looking ahead, we envision AR-driven approaches playing a central role in shaping agile standards that evolve alongside the technologies they govern.
Acknowledgements. We thank Helen Oakley and Rhea Michael Anthony for the industrial validation in Section 5.4; Prof. Daniel M. German and Prof. Zhen Ming (Jack) Jiang for their mentoring; and the members of the SPDX AI and Dataset Working Group for their contributions and feedback. Arthit Suriyawongkul’s contribution to the WG is made possible through the financial support of Taighde Éireann – Research Ireland under Grant number 18/CRT/6224 (Research Ireland Centre for Research Training in Digitally-Enhanced Reality (d-real)).